Calculate lambda for the following table?

Catholic Protestant Jewish Other
Male 6 10 16 17
Female 14 10 8 6
Answer #1

\lambda ={\frac {\varepsilon _{1}-\varepsilon _{2}}{\varepsilon _{1}}}.

\lambda = \frac{49 - (6 + 14)}{49} = 0.592